If the same amount of energy in the form of heat is added to 5.00 g samples of metals Ag, Al, Cu, which metal will undergo the largest temperature change? The heat capacities of the metals in J/g 0 C are: 0.235, 0.897, 0.385 respectively.


1
Expert's answer
2022-02-17T04:36:02-0500

"Q= m\u00d7 c\u00d7 dT"

If the amount of energy =500J

Then;


Change in temperature for Ag is

500= 5× 0.235× dT

dT=425.5°c

Change in temperature for Al

500= 5× 0.897× dT

dT= 111.5°C

Change in temperature for Cu

500= 5× 0.385× dT

dT= 259.7°c


Ag Will undergo the largest temperature change